# MotherNet: A Foundational Hypernetwork for Tabular Classification

### v1

@@ Abstract
The advent of Foundation Models is transforming machine learning across many modalities (e.g., language, images, videos) with prompt engineering replacing training in many settings. Recent work on tabular data (e.g., TabPFN) hints at a similar opportunity to build Foundation Models for classification for numerical data. In this paper, we go one step further and propose a hypernetwork architecture that we call MotherNet, trained on millions of classification tasks, that, once prompted with a never-seen-before training set generates the weights of a trained “child” neural-network. Like other Foundation Models, MotherNet replaces training on specific datasets with in-context learning through a single forward pass. In contrast to existing hypernetworks that were either task-specific or trained for relatively constraint multi-task settings, MotherNet is trained to generate networks to perform multiclass classification on arbitrary tabular datasets without any dataset specific gradient descent.
The child network generated by MotherNet using in-context learning outperforms neural networks trained using gradient descent on small datasets, and is competitive with predictions by TabPFN and standard ML methods like Gradient Boosting. Unlike a direct application of transformer models like TabPFN, MotherNet generated networks are highly efficient at inference time. This methodology opens up a new approach to building predictive models on tabular data that is both efficient and robust, without any dataset-specific training.
Machine Learning, ICML
[…]
@@ 1 Introduction
Foundation Models, i.e., large transformer-based (Vaswani et al., 2017) models trained on massive corpora, are disrupting machine learning in many areas such as natural language and reasoning tasks. The field is quickly moving from small task-specific models to large generic models with task-specific instructions via prompting and in-context learning.
This tidal wave of change has yet to reach tabular data, the most common data type in real-world machine learning applications (Chui et al., 2018), where currently traditional machine learning methods are still the norm. In this paper, we explore a new angle of applying transformer-based Foundational Models to tabular classification. Such models have the potential to transform ML for Tabular data, replacing costly and slow AutoML activities with in-context learning. The existing TabPFN approache, while very promising in terms of accuracy, fall short of state-of-the-art classical solutions in terms of scalability to large training datasets and inference runtime. We focus our attention towards removing these limitations.
@@
We introduce a new architecture, called MotherNet, which adapts the TabPFN transformer architecture (Hollmann et al., 2022; Vaswani et al., 2017) to produce model weights for a neural network of fixed structure (a two hidden layer neural network in the experiments ) that performs competitively with baseline methods such as gradient boosting (Friedman, 2001; Chen & Guestrin, 2016). Using MotherNet to generate neural network weights substantially outperforms learning by gradient descent in computational efficiency, accuracy and ease of use on small tabular datasets.
@@
Our approach combines the transformer architecture of TabPFN (Hollmann et al., 2022) with the idea of hyper-networks (Ha et al., 2017), to produce state-of-the-art classification models in a single forward pass. Unlike original work in hypernetworks (Ha et al., 2017), which used a small hyper network to generate a large “main” network, we are training a large, transformer-style hyper-network to generate a compact classification network. Compared to the approach of Hollmann et al. (2022), this allows for very fast inference, making the model applicable for low-latency requirements and for prediction on large datasets. Compared to earlier work on hypernetworks, we train a single hyper-network to address tabular classification on numeric data in general, i.e. in style of a foundational model, instead of a task-specific or multi-task hypernetwork.
@@
Our key contribution is demonstrating that it is indeed possible to generate neural networks directly as the output of a transformer model, without the need to do any dataset-specific learning or gradient descent. Using a fixed model structure, we are able to produce neural networks that work well on small numeric tabular datasets from the OpenML CC-18 benchmark suite (Bischl et al., 2017).
The key intuition is that the “mother” network has been trained on millions of classification scenarios and has learned a certain degree of regularization. This “wisdom” is passed down to the child, allowing for the child to operate in a label-scarce environment without over-reacting to outliers present in the prompt.
Aside from applications to AutoML, our approach also has an interesting theoretical angle. Given a tabular training dataset, MotherNet learns (via in-context learning) to build a machine learning model in the form of an MLP. However, throughout the procedure, there is no optimization for training set accuracy at any point. Therefore, the MotherNet approach does not use framework of regularized empirical risk minimization that is the basis of nearly all modern machine learning methods. Instead, it directly optimizes expected test-set performance, with distributional assumptions encoded in the training of the MotherNet model.
Our findings can be summarized as follows:
  • It is possible to create accurate small neural networks without gradient descent, using in-context learning with a transformer.
  • Our approach performs competitively with baseline methods, including Gradient Boosting with hyper-parameter tuning, without any dataset specific tuning of our method.
  • Our approach has extremely fast inference time, as well as combined inference and training time.
  • Our approach provides a proof-of-concept for building standard machine learning models without relying on empirical risk minimization, therefore providing a new perspective on generalization.
Figure 1: MotherNet architecture. Given training data $(x_{1},y_{1}),\dots,(x_{r},y_{r})$, the transformer produces a vector $\phi$, which is reshaped to weight matrices of an MLP with low-rank weight structure.
@@ 2 Related Work
2.1 Foundation Models
The use of Foundation Models for natural language tasks has taken research and applications by storm, especially since the launch of ChatGPT and GPT-4 (OpenAI, 2023). One of the surprising observations about the work in this area is that models trained purely for next word prediction can be adapted to novel NLP tasks by providing the correct prompt (Brown et al., 2020). This approach replaces training or even fine-tuning by providing task-dependent input to a fixed model, also known as in-context learning. While the original training of the model might be quite expensive (computationally, data needed, and financially), once the model is built, in-context learning allows for near-instantaneous adaptation to new tasks and new contexts. Much work has been done to further improve the responsiveness to prompts, and adjust output to human preference. Our paper addresses tabular numeric data, and in this context, providing task definitions as part of the input is non-trivial. Therefore we fix the task to multi-class classification. Our goal is to produce a Foundation Model that can generalize this task across training datasets, in contrast to state-of-the-art approaches that train a classifier per dataset.
2.2 Transformers for Tabular Data
There have been several works investigating the use of small, heterogeneous tables for question answering, and extracting tables from data, using large language models and specifically fine-tuned transformers (Yin et al., 2020; Iida et al., 2021).
While generic language models can be adapted to work on tables, the restriction on the number of input tokens, even for the largest models, severely limits the amount of data that can be ingested. Using a dedicated architecture for numeric data allows scaling to much larger datasets. More specifically, compare the memory usage of a large language model like LLAMA (Touvron et al., 2023)111we choose LLAMA over the GPT family since the tokenization used in GPT makes it less suitable for dealing with numeric data with that of a transformer for numeric data such as TabPFN (Hollmann et al., 2022). LLAMA has a vocabulary size of 32,000 and an embedding dimension of 4096 (in the smallest version). Expressing a single feature in the iris dataset, which are decimals with two places, takes four tokens, so approximately $4096\cdot 4=16384$ floating point numbers or a sparse vector of length $32,000\cdot 4=128000$. Using a transformer for floating point tables, it takes a single floating point number. Expressing the whole iris dataset (four features, three classes, 150 data points) in a way consumable by a large language model yields 5000 to 15000 tokens, depending on the tokenizer and JSON representation, which is above the capability of most large language models: Llama supports 2048, the largest current model, GPT-4, up to 32k, while Hollmann et al. (2022) showed scaling to 100 features and 1000 data points (up to 5000 in extrapolation experiments). Notably, the token limits on the Large Language Model combine input, output and instructions, while the limitations in Hollmann et al. (2022) are for the training set, with the assumption of a test set of the same size.
@@ 2.3 TabPFN
Recently Hollmann et al. (2022), building on the work of Müller et al. (2021), introduced a transformer architecture that is capable of performing supervised classification on tabular numeric data. This work is quite distinct from other transformer architectures on tabular data in that it is focused on numeric input and numeric output. The authors design a prior over synthetically generated datasets, based on Structural Causal models and Bayesian Neural Networks. Using draws from this prior, they are able to train a model that generalizes to perform supervised classification on real-world tabular datasets. Our work builds on the work of Hollmann et al. (2022), and adds the capability to create a dataset-specific model. This allows much faster and more memory-efficient inference given a fixed training set, and allows standard model inspection techniques for neural networks, as opposed to inspecting the activations of the transformer model. We use the prior of Müller et al. (2021), therefore sharing their assumptions about the data distribution, and as we are using a transformer architecture, we are also limited in the training set size similar to their approach.
@@ 2.4 Learning to Learn and Meta-learning
There has been a long history of approaches to “learning to learn” and to build neural network that produce other neural networks in machine learning (Schmidhuber, 1992; Ravi & Larochelle, 2016; Andrychowicz et al., 2016; Thrun & Pratt, 1998). Given the long history, we will only review some more recent and closely related approaches.
Ha et al. (2017) introduce the term hypernetwork for networks that produce networks using task-specific embeddings. The hypernetwork and embedding are both learned via gradient descent on the same dataset, reducing the approach (in the non-recurrent case) in essence to a standard neural network with a low-rank structure in the weights.
Later, Finn et al. (2017) proposed the MAML framework for multi-task learning, which generates “easy-to-finetune” weights for a fixed architecture, which allows quick adaption to new tasks. Among the many follow-ups to MAML, of particular relevance to this work is CAVIA (Zintgraf et al., 2019), which decomposes the network into a task-specific part to be fine-tuned in the ”inner loop” and a task agnostic part, which is fixed after meta-learning. This approach has been studied more in-depth, and put into the context of hypernetworks in Zhao et al. (2020). Zhao et al. (2020) learn an initialization for the task embedding using a MAML-style objective, which can then be further fine-tuned using gradient descent.
Similar to our work, MAML and later variants aim to speed up and improve learning of a network by meta-learning from a related set of tasks, using a fixed network architecture. However, MAML requires performing gradient descent to adjust to specific tasks, and factorization based approaches like Zintgraf et al. (2019) and Zhao et al. (2020) still require gradient descent on the task-specific components of the model, while MotherNet produces task-specific weights using a HyperNetwork, requiring no further gradient descent. Furthermore, MAML and related multi-task methods have usually been applied to tasks that require large models, in particular computer vision tasks, with close relationships between tasks, i.e. learning new classes in an image classification dataset.
Not relying on gradient descent not only provides improved computational performance, but it fundamentally changes the nature of the optimization. In CAVIA and MAML-style approaches, the network weights are (directly or indirectly) optimized to minimize training-set loss on a new task, while MotherNet never optimizes training-set loss, and so does not suffer from overfitting (though potentially from meta-overfitting to the meta-trainingset). This removes the need for any regularization for generating the task-specific model.
While it would be possible to adapt MAML to the setting of variable number of features and classes by borrowing methods from Hollmann et al. (2022), this is beyond the scope of this work.
More closely aligned with our work, Bertinetto et al. (2016) also propose a gradient-free approach to produce fully-fledge student networks based on single-shot examples in OCR and object tracking. Their objective and formulation closely resembles ours; however, in this work, a single “exemplar” is a tabular training dataset, while in Bertinetto et al. (2016), it is a single handwritten digit, or example of an object to track. By learning to ingest generic tabular datasets, our approach successfully deals with a much broader variety of tasks.
In spirit, our approach is also related to more classical meta-learning approaches, that have attempted at generalizing knowledge among datasets  (Vanschoren, 2018). For tabular data, this has usually been in the form of hyper-parameters, either via the generation of portfolios (Brazdil et al., 2003; Wistuba et al., 2015; Abdulrahman et al., 2018) or by using dataset characteristics such as meta-features (Rivolli et al., 2018). Our work differs, in that our meta model takes in the whole dataset, and, instead of producing hyper-parameters or candidate model architectures, it directly produces a fully trained model. This makes the MotherNet architecture much cheaper than traditional meta-learning approaches, in that no actual per-dataset model training needs to be performed in the meta-training or prediction phases. Commonly, meta-learning also requires local search over hyper-parameters, in addition to the use of a meta-model (Hutter et al., 2019), while we found our one-shot weights to work robustly without any modifications.
@@ 3 Methodology
3.1 Background
Hollmann et al. (2022) introduced TabPFN, an adaption of the transformer architecture to solve tabular classification problems. TabPFN uses a transformer where each input “token” is a row of the tabular dataset. The model is adapted to work with a variable number of features by zero-padding (and scaling) to 100 features. For the training data, linear projections of the input rows are summed with linear projections of integer classification labels. For the test data, the addition of labels is simply omitted, and class-probabilities are produced as output tokens. The model is trained to minimize cross-entropy of probabilistic predictions on the test data points. Attention is masked so that all training points can attend to all other training points, while test points can only attend to training points. A variable number of classes is handled by training for up to ten classes, and when predicting for a dataset with $k\leq 10$ classes, using only the first $k$ outputs in the softmax layer.
TabPFN showed strong predictive performance without any per-dataset tuning, and with extremely fast time to train and predict on small datasets ($\leq 2000$ data points). Because of the quadratic nature of the self-attention matrix, training on larger datasets is impractical with the method proposed in Hollmann et al. (2022). Their prior focused on numeric-only data, so their model was evaluated on datasets without categorical or missing features. We extend this approach in our work removing some of the limitations discussed next.
@@ 3.1 Background
Comparing speed and computational efficiency between TabPFN and traditional ML and AutoML methods is somewhat complicated, as they have very different characteristics. In particular, there is no dataset specific training phase after meta-training when applying TabPFN, only near-instantaneous in-context learning. Prediction, on the other hand, is significantly slower than prediction in standard ML models, as predicting on $N_{p}$ data points given a training dataset of size $N_{t}$ requires constructing attention matrices of sizes $N_{t}\times N_{t}$ and $N_{t}\times N_{p}$. This is similar to the characteristics of traditional $kNN$ classifiers or Gaussian Process models, which (absent specific optimizing data structures) have no training phase.
The gold standard of modern tabular learning, gradient boosted trees (Friedman, 2001; Chen & Guestrin, 2016; Ke et al., 2017), on the other hand, has significant training cost, in particular when accounting for hyper-parameter tuning, but has prediction cost independent of the training set size, and only dependent on the model size.
From these, it is clear that applying TabPFN to large training sets using the standard transformer architecture is challenging. Moreover, when looking at just prediction time, TabPFN is far behind gradient boosted models.
While both models have prediction time linear in the number of test data points $N_{t}$, the constant in the case of TabPFN is much larger. Together with the memory requirements of a large transformer model, this makes TabPFN impractical for settings where fast, on-demand predictions are required.
Next we present two approaches, a surprisingly effective baseline distillation approach, and MotherNet, that address some of the runtime and scalability limitations of TabPFN.
Figure 2: Comparing decision boundaries on synthetic toy datasets, adapted from the scikit-learn (Pedregosa et al., 2011) documentation. MotherNet decision boundaries closely resemble TabPFN and traditional Neural Network boundaries.
3.2 Distillation
While TabPFN has good training-time properties, it has high prediction time cost. A natural solution is to create a dataset specific compression or instantiation of the model based on a given training dataset. There are many potential routes to this, including model pruning or distilling the transformer model. We chose to evaluate a simple baseline, applying the predictions of TabPFN as a teacher model for a small feed-forward neural network, such as a Multilayer Perceptron (MLP) model, that is trained specifically for a dataset, analogous to the methodology of Hinton et al. (2015). Note that we are not attempting to distill TabPFN, but instead create dataset-specific distillations for each dataset we want to predict on. Furthermore, while TabPFN is acting as a teacher model, it never saw the dataset for which it is a teacher during training time. While this is a natural way to distill the model for dataset-specific prediction, we are not aware of this being investigated before. The reduction in model size and complexity between the original model and the distilled model is dramatic (at nearly 3 orders of magnitude): We use an MLP model with two hidden layers of size 128, which results in a maximum of  30k parameters (with 100 input features and 10 targets, i.e., the limit in the datasets we consider as most datasets are smaller), while the original TabPFN has  26M parameters.
In Section 4.2 we show that this distillation process is extremely effective in approximating the predictions made by by the original model without any dataset-specific tuning of the MLP hyper-parameters. This is in stark contrast to training the same MLP from scratch using gradient descent, which is unable to achieve competitive results without hyper-parameter tuning. Creating the distilled model introduces a distinct training phase in using TabPFN, but tremendously accelerates prediction.
Next we present a novel approach that also produces a small ”child” network, without incurring the cost of a distinct training phase.

### v2

# MotherNet: Fast Training and Inference via Hyper-Network Transformers

@@ Abstract
Foundation models are transforming machine learning across many modalities, with in-context learning replacing classical model training. Recent work on tabular data hints at a similar opportunity to build foundation models for classification for numerical data. However, existing meta-learning approaches can not compete with tree-based methods in terms of inference time. In this paper, we propose MotherNet, a hypernetwork architecture trained on synthetic classification tasks that, once prompted with a never-seen-before training set generates the weights of a trained “child” neural-network by in-context learning using a single forward pass. In contrast to most existing hypernetworks that are usually trained for relatively constrained multi-task settings, MotherNet can create models for multiclass classification on arbitrary tabular datasets without any dataset specific gradient descent. The child network generated by MotherNet outperforms neural networks trained using gradient descent on small datasets, and is comparable to predictions by TabPFN and standard ML methods like Gradient Boosting. Unlike a direct application of TabPFN, MotherNet generated networks are highly efficient at inference time. We also demonstrate that HyperFast is unable to perform effective in-context learning on small datasets, and heavily relies on dataset specific fine-tuning and hyper-parameter tuning, while MotherNet requires no fine-tuning or per-dataset hyper-parameters.
@@ 1 Introduction
Foundation models, i.e., large transformer-based (Vaswani et al., 2017) models trained on massive corpora, are disrupting machine learning in many areas such as natural language and reasoning tasks. These domains shifted from small task-specific models to large generic models with task-specific instructions via prompting and in-context learning. However, this shift has not yet reached tabular data, the most common data type in real-world machine learning applications (Chui et al., 2018), which is still dominated by traditional machine learning methods and neural networks with in-weight learning. This paper explores a new approach to applying transformer-based Foundational Models to tabular classification, demonstrating their potential to replace costly and slow AutoML with in-context learning. The existing TabPFN (Hollmann et al., 2022) approach is promising in terms of accuracy and training time, but falls short of state-of-the-art classical solutions in terms of training set scale, being restricted to $1000$ to $3000$ data points for training, and inference runtime, being approximately ten times slower to predict than a comparable tree-based method.
@@
We introduce a new architecture, called MotherNet, which adapts the TabPFN architecture to a hypernetwork setup to produce model weights for a feed forward neural network. Our method performs competitively with baseline methods such as gradient boosting (Friedman, 2001; Chen & Guestrin, 2016) and outperforms learning a neural network with gradient descent, being much faster to train, providing higher accuracy and requiring no tuning of hyperparameters on small tabular datasets. Our approach combines the transformer architecture of TabPFN (Hollmann et al., 2022) with the idea of hypernetworks (Ha et al., 2017), to produce state-of-the-art classification models in a single forward pass. Unlike original work in hypernetworks (Ha et al., 2017), which used a small hyper network to generate a large “main” network, we are training a large, transformer-based hypernetwork to generate a compact classification network.
@@
Compared to the approach of Hollmann et al. (2022), this allows for much shorter prediction time. However, just as TabPFN, MotherNet is restricted by the quadratic memory requirements of the transformer architecture, and does not scale well above approximately 5,000 data points. Compared to earlier work on hypernetworks, we train a single hypernetwork to address tabular classification on numeric data in general, i.e. in the style of a foundational model, instead of a task-specific or multi-task hypernetwork.
@@
We demonstrate that it is possible to generate neural networks directly as the output of a transformer model, without the need to do any dataset-specific learning or gradient descent. Using a fixed model structure, we are able to produce neural networks that work well on small numeric tabular datasets from the OpenML CC-18 benchmark suite (Bischl et al., 2017), and show that our approach also provides a good trade-off of speed and accuracy on the TabZilla dataset collection McElfresh et al. (2024). Training and inference code and pre-trained model weights are made publicly available  [^1] https://github.com/microsoft/ticl.
@@ 2 Related Work
2.1 LLMs for Tabular Data
There have been several works investigating the use of small, heterogeneous tables for question answering, and extracting tables from data, using large language models (LLMs) and specifically fine-tuned transformers (Yin et al., 2020; Iida et al., 2021). These architectures are table-specific, but meant to answer natural language questions. Our work focuses on supervised classification. Separately approaches have been proposed to apply Large Language Models directly to supervised tabular classification tasks (Dinh et al., 2022; Hegselmann et al., 2023). These works generally require tokenization on the level of each input feature. This allows including world-knowledge about feature names and potentially feature values, but comes at an extreme computational cost, and strong limitations on the number of features and data points, as the size of the attention matrix is quadratic in both features and samples, and has a non-trivial constant factor for encoding floating point numbers into separated string values.
@@ 2.2 TabPFN
Recently Hollmann et al. (2022), building on the work of Müller et al. (2021), introduced a transformer architecture that is capable of performing supervised classification on tabular numeric data. This work is quite distinct from other transformer architectures on tabular data in that it is focused on numeric input and numeric output.
TabPFN uses a transformer where each input “token” is a row of the tabular dataset. The model is adapted to work with a variable number of features by zero-padding (and scaling) to 100 features. For the training data, linear projections of the input rows are summed with linear projections of integer classification labels. For the test data, the labels are omitted and class-probabilities are produced as output tokens. The model is trained to minimize cross-entropy on the test data points. Attention is masked so that all training points can attend to all other training points, while test points can only attend to training points. A variable number of classes is handled by training for up to ten classes, and when predicting for a dataset with $k\leq 10$ classes, using only the first $k$ outputs in the softmax layer. The authors design a prior over synthetically generated datasets, based on Structural Causal models and Bayesian Neural Networks. Using draws from this prior, they are able to train a model that generalizes to perform supervised classification on real-world tabular datasets. TabPFN showed strong predictive performance without any per-dataset tuning, and with extremely fast time to train and predict on small datasets ($\leq 3000$ data points) (McElfresh et al., 2024). Because of the quadratic nature of the self-attention matrix, training on larger datasets is impractical with the method proposed in Hollmann et al. (2022). Our work builds on the work of Hollmann et al. (2022), and adds the capability to create a dataset-specific model.
@@ 2.3 Learning to Learn and Meta-learning
There has been a long history of approaches to “learning to learn” and to build neural network that produce other neural networks (Schmidhuber, 1992; Ravi & Larochelle, 2016; Andrychowicz et al., 2016; Thrun & Pratt, 1998). Given the long history, we will only review some more recent and closely related approaches. Most approaches solve transfer-learning or task-adoption within a fairly narrow family of tasks, often one-dimensional regression or character recognition, while our work addresses classification on any small tabular dataset. Ha et al. (2017) introduce the term hypernetwork for networks that produce networks using task-specific embeddings. The hypernetwork and embedding are both learned via gradient descent on the same dataset, reducing the approach (in the non-recurrent case) to a standard neural network with a low-rank structure in its weights. Bertinetto et al. (2016) propose a gradient-free approach to produce student networks based on single-shot examples in OCR and object tracking. Their objective and formulation closely resembles ours; however, in this work, a single “exemplar” is a tabular training dataset, while in Bertinetto et al. (2016), it is a single handwritten digit, or a single object to track. A transformer based approach for a hypernetwork generating convolutional neural networks has been investigated in Zhmoginov et al. (2022). Conditional Neural Processes (Garnelo et al., 2018) also perform task adaption without gradient descent. In contrast to the original work of Garnelo et al. (2018), we are using a transformer architecture instead of a feed-forward neural network, and are able to address a much wider range of tasks. While later work on Neural Processes used transformers (Nguyen & Grover, 2022) in an architecture closely related to TabPFN, we are not aware of an implementation of Conditional Neural Processes using transformers, which would yield an architecture more similar to MotherNet. Most recently, Bonet et al. (2023) introduced HyperFast, a hypernetwork that, similar to MotherNet, is trained to perform classification on generic tabular datasets. HyperFast avoids the quadratic complexity of the transformer attention mechanism, and therefore scales to larger datasets. We compare to HyperFast with and without gradient descent and with hyper-parameter tuning in Section 4. and find that despite the large overlap of the HyperFast training set and our test-set, HyperFast without gradient descent performs poorly, and hyper-parameter tuning is necessary for good performance.
@@ 3 Methodology
Hollmann et al. (2022) introduced TabPFN, an adaption of the transformer architecture to solve tabular classification problems. As this work closely builds on TabPFN, we want to briefly review its architecture. TabPFN uses a transformer where each input “token” is a row of the tabular dataset. The model is adapted to work with a variable number of features by zero-padding (and scaling) to 100 features. For the training data, linear projections of the input rows are summed with linear projections of integer classification labels. For the test data, the labels are omitted and class-probabilities are produced as output tokens. The model is trained to minimize cross-entropy on the test data points. Attention is masked so that all training points can attend to all other training points, while test points can only attend to training points. A variable number of classes is handled by training for up to ten classes, and when predicting for a dataset with $k\leq 10$ classes, using only the first $k$ outputs in the softmax layer. TabPFN showed strong predictive performance without any per-dataset tuning, and with extremely fast time to train and predict on small datasets ($\leq 3000$ data points) (McElfresh et al., 2024). Because of the quadratic nature of the self-attention matrix, training on larger datasets is impractical with the method proposed in Hollmann et al. (2022).
@@
Comparing speed and computational efficiency between TabPFN and traditional ML and AutoML methods is somewhat complicated, as they have very different characteristics. In particular, there is no dataset specific training phase after meta-training when applying TabPFN, only near-instantaneous in-context learning. Prediction, on the other hand, is significantly slower than prediction in standard ML models, as prediction requires computing attention between test and training data. On the other hand, gradient boosted trees (Friedman, 2001; Chen & Guestrin, 2016; Ke et al., 2017) have significant training cost, in particular when accounting for hyper-parameter tuning, but are extremely fast for prediction. Together with the memory requirements of a large transformer model, this makes TabPFN impractical for settings where fast, on-demand predictions are required. Next we present two approaches, an effective baseline distillation approach, and MotherNet, that address some of the runtime and scalability limitations of TabPFN.
Figure 1: MotherNet architecture. Given training data $(x_{1},y_{1}),\dots,(x_{r},y_{r})$, the transformer produces a vector $\phi$, which is reshaped to weight matrices of an MLP with low-rank weight structure. Green blocks are individual data points and their activations, orange layers are activations created during in-context learning or in the forward-pass during meta-learning, and light blue layers are learned during meta-training.